### v4

# BlueME: Robust Underwater Robot-to-Robot Communication Using Compact Magnetoelectric Antennas

Abstract
We present the design, development, and experimental validation of BlueME, a compact magnetoelectric (ME) antenna array system for underwater robot-to-robot communication. BlueME employs ME antennas operating at their natural mechanical resonance frequency to efficiently transmit and receive very-low and low-frequency (VLF/LF) electromagnetic signals underwater. We detail the system’s design, simulation, fabrication, and integration onto low-power embedded platforms, emphasizing portability and scalability. For performance evaluation, we deployed BlueME on mobile robot platforms in freshwater (lake) and saltwater (ocean) trials. Our tests demonstrate reliable signal transmission and detection using the BlueME antenna system at distances beyond 700 meters while consuming less than 10 watts of power. Ocean trials demonstrate that the system operates effectively under challenging conditions, such as turbidity, obstacles, and multipath interference, which are factors that typically degrade acoustic and optical methods. Our analysis also examines the impact of complete submersion on system performance and identifies key deployment considerations. This work represents the first practical underwater deployment of ME antennas outside the laboratory and implements the largest VLF/LF ME array system to date. BlueME demonstrates significant potential for marine robotics and automation in multi-robot cooperative systems and remote sensor networks.
Keywords. Marine Robotics; Underwater Communication; Magnetoelectric Antennas.
1 Introduction
Underwater robot-to-robot communication is essential for ensuring the coordinated operation of autonomous underwater vehicles (AUVs) in a wide range of applications including environmental monitoring, search and rescue, and scientific expeditions [1, 2]. In subsea environments where human intervention is limited, AUVs rely on seamless communication to share information, synchronize tasks, and respond to dynamic changes in real time [3, 4, 5] in multi-robot missions or when communicating with surface vessels [6, 7, 8]. This capability is particularly crucial in surveying or mapping deep-sea areas [4], navigating overhead structures [9], and collecting samples from areas inaccessible to human divers [7].
Conventional electromagnetic and radio frequency (RF) signals are severely attenuated underwater [1, 10], limiting their effective communication range to only a few meters. While propagation improves at very-low and low-frequencies (VLFs/LFs) [11], it typically comes at the cost of high transmission power and impractically large antenna sizes [12, 13]. In seawater, electromagnetic signal attenuation ranges from approximately 1 to 10 dB per meter depending on both frequency and conductivity [14]. This high attenuation severely limits the viability of RF communication for marine robotics applications [15].
Acoustic communication has been the most commonly used modality for underwater robot communications, offering omnidirectional transmission and acceptable signal attenuation over longer distances [16, 17]. Long-range low-frequency systems are used in ocean-based communication [18], while short-range high-frequency systems are preferred for robot positioning and data transfer [19, 20]. Acoustic communications suffer from Doppler effects, phase and magnitude fluctuations, and multi-path interference [21, 22, 23], compromising reliability and efficiency. Acoustic signals can be disruptive to marine ecosystems as well, limiting their long-term deployments. Optical communication systems (e.g., visible light, blue-green lasers) offer higher data rates [24, 25, 26], but are limited by line-of-sight requirements. They are susceptible to biofouling [27], a process where microorganisms and biological buildup degrade optical sensors and signal quality over time [28].
Magnetoelectric (ME) antennas offer a promising solution to overcome the limitations of acoustic and optical communication in underwater environments [29, 11, 30, 31, 32]. By leveraging the magnetoelectric effect, these antennas efficiently transmit and receive electromagnetic signals with a more compact design than traditional electrically small antennas (ESAs) [33]. This efficiency, combined with their resilience to multi-path interference, non-line-of-sight requirements, and Doppler effects, makes ME antennas well-suited for space- and power-constrained underwater applications.
In this paper, we present the design, development, and validation of BlueME, a novel magnetoelectric (ME) antenna system for real-time data transmission between underwater robots. As shown in Fig. 1, we use a compact array of ME antennas integrated into pressure-compensated enclosures for real-time data transmission and retrieval between robots and sensors deployed underwater. While previous ME-based underwater communication efforts have primarily focused on basic connectivity tests in controlled water-tank environments [29, 34, 35, 11], we validate BlueME in a natural lake and ocean, demonstrating its potential as a deployable communication system for practical marine robotics applications.
In particular, we deploy the proposed system on ASVs (Autonomous Surface Vehicles) and underwater ROVs (Remotely Operated Vehicles) for comprehensive performance evaluation and feasibility analysis. Our field experimental results demonstrate that the BlueME system offers a more robust, long-range, scalable, and efficient alternative to the traditional acoustic and optical communication modalities [36, 37]. We also analyze the antenna characteristics, underlying theory of operations, and relevant engineering constructs for practical deployments of ME antenna arrays underwater.
Overall, we make the following contributions in this work:
We design, simulate, and fabricate an ME antenna array system (BlueME) for real-time underwater communication between mobile robots and/or sensor nodes. This is the first ME antenna array system tested in practical environments beyond controlled water-tank setups.
We present a seamless integration of BlueME into pressure-compensated enclosures to ensure reliable underwater operations in real-time. We demonstrate that the enclosed system is compact and end-to-end portable for use on low-power embedded devices.
We validate the system through simulation as well as field experimental trials in freshwater (lake) and saltwater (ocean) environments, demonstrating its effectiveness for underwater communication. The system performs reliably when fully submerged and is unaffected by turbidity, line-of-sight obstacles, and shallow-water interference, which limit acoustic and optical systems.
Our field deployments reveal that we can achieve a high-fidelity communication link between mobile robots for up to meters with a power footprint of only - watt. These capabilities demonstrate its potential for robust and low-latency communication underwater.
2 Background & Related Work
2.1 Underwater Communication Technologies
Underwater communication is crucial for data exchange among autonomous underwater vehicles (AUVs), submarines, and subsea sensor networks [10, 16, 38]. The most prevalent technology is acoustic communication, operating within the 10 Hz to 100 kHz frequency bands and supporting data rates up to several hundred Kbps. Despite its widespread use, acoustic communication faces significant limitations, including variable propagation delays and Doppler shifts due to the relatively slow speed of sound underwater (approximately 1500 m/s) [22, 39]. Phase and magnitude fluctuations, along with multi-path interference, further degrade signal reliability[13, 40]. Additionally, acoustic signals can adversely affect marine life and are susceptible to environmental noise [41, 42].
Optical communication offers higher data rates and lower latency compared to acoustic methods [24, 25]. Notable systems include the BlueComm system, achieving 20 Mbps over distances up to 200 m [43]; the Ambalux system, providing 10 Mbps over 40 m [25]; and LUMA modems, supporting 10 Mbps across distances up to 50 m [44]. Underwater visible light communication (UVLC) has also been explored for extended ranges using wavelengths between 450–550 nm, which experience lower attenuation in water [26, 28]. However, optics are limited by line-of-sight requirements and are susceptible to scattering, absorption due to turbidity, and biofouling effects that degrade performance over time [45, 46].
Radio frequency (RF) electromagnetic communication underwater has historically been challenging due to high attenuation of EM waves in conductive media like seawater [14]. Attenuation increases with frequency, limiting effective communication to short distances [15]. However, RF communication offers advantages such as the ability to cross air-water and water-sediment boundaries, operation in turbid conditions, and non-line-of-sight capabilities [10, 47]. Recent research has revisited underwater RF communication using very low frequency (VLF, 3–30 kHz) and ultra-low frequency (ULF, 300–3000 Hz) bands to mitigate attenuation [29, 10]. Despite these advances, practical deployment is hindered by the need for large antennas and/or high transmission power, making it less suitable for mobile underwater applications [48, 12].
2.2 ME Antennas for Low-Power Communication
ME antennas have emerged as a prospective solution for underwater communication at low frequencies with reduced antenna sizes [11]. ME antenna structures utilize the magnetoelectric effect, where a magnetic field induces strain in a magnetostrictive material, which is transferred to a piezoelectric material to generate an electric field, and vice versa [49, 50]; see Fig. 2. This mechanism enables efficient energy conversion in compact configurations, addressing the size limitations of traditional electrically small antennas (ESAs).
Research on ME communication systems has primarily focused on biomedical applications such as implantable devices [51, 52]. More recently, studies have explored their potential for underwater applications [29, 34, 35]. Fabrication typically involves layering magnetostrictive materials like Metglas with piezoelectric materials (e.g., PZT, PVDF), bonded using structural epoxy, and integrating conductive elements [53]. ME antennas operating in VLF range can overcome some limitations of traditional RF communication by providing efficient transmission without requiring impractically large antennas. They are less susceptible to multi-path interference and can function in non-line-of-sight conditions.
However, significant challenges remain in designing and fabricating compact, low-power ME antennas suitable for integration with mobile robotic platforms [54, 55]. Key issues include optimizing antenna performance, developing feasible deployment strategies, and characterizing communication properties in dynamic underwater conditions. This work aims to advance the field by developing a novel ME antenna array system tailored for marine robotics applications, addressing current gaps in practical deployment and performance evaluation.
3 BlueME System Design
3.1 ME Antenna Design: Theory and Simulation
We first investigate the fundamental mechanical resonance frequency of the ME antenna for underwater communication. Considering the requirements of compact size, low attenuation characteristics, and long propagation distances, we calculate a frequency of approximately kHz as follows [56]:
Here, = mm is the sample length; = GPa and = GPa are the Young’s moduli [56]; = and = are the volume fractions; = Kg/m3 and = Kg/m3 are the densities of PZT and Metglas, respectively [57, 58]. In our design, PZT-5J is chosen as the piezoelectric layer, and Metglas as the magnetostrictive layer.
Fig. 2 illustrates an outline of our design showing how magnetostriction and piezoelectricity are coupled together in ME antennas. The ME effect is characterized by the magnetoelectric coupling coefficient , defined as the induced electric field per unit applied magnetic field [59, 60]:
where is the electric field, is the magnetic field, is the voltage generated across the piezoelectric layer, and is its thickness. The ME coupling arises from the product of the magnetostrictive and piezoelectric effects, and our objective is to maximize the strain transfer between the magnetostrictive and piezoelectric layers to increase .
To validate this theoretical analysis, we use COMSOL Multiphysics [61] to simulate a three-layer ME structure. As shown in Fig. 3, we evaluate the displacement magnitude of our antenna with a Metglas-PZT-Metglas thickness of -- m. By simulation, we predict the eigenfrequency for the bias-free fundamental displacement mode to be kHz.
Antenna Enclosure and Array Assembly. As shown in Fig. 4, each ME antenna is housed in a custom 3D-printed enclosure. The upper shell of the enclosure contains a pogo pin and two vertical AWG copper wires; the pogo pin is bonded to a designated antenna hole using a UV-curable cyanoacrylate adhesive. The AWG wires from the antenna are twisted and soldered to the AWG wires, which extend through the lower shell. The two shells are then enclosed and arranged in a grid within a frame that includes twenty N52 neodymium magnets. Each magnet is positioned from the antenna edge, all oriented with the same north–south polarity. The magnetic bias distribution across a representative antenna was measured as shown earlier in Fig. 3. Finally, the receiver and transmitter antennas are connected in series and parallel configurations, respectively.
3.2 Pressure Compensation & System Integration
The antenna arrays and associated wiring are enclosed within a lidded resin-printed shell; see Fig. 4. A laser-cut silicone gasket ensures a watertight seal, secured by sixteen stainless steel screws torqued to inch-lbs. The enclosure is then filled with a pressure-compensating oil to mitigate hydrostatic pressure effects.
In our experiments, a transmitter antenna array is mounted on an autonomous surface vehicle (ASV) (BlueRobotics BlueBoat), while a receiver array is mounted on an underwater remotely operated vehicle (ROV) (BlueRobotics BlueROV2). As shown in Fig. 5, the transmitter array connects to a power amplifier driven by a signal from a Digilent Analog Discovery 2 (AD2) board.
The receiver array connects to a low-noise amplifier (AlphaLab LNA10) with configured gain and filtering, which is interfaced with a Digilent Analog Discovery 3 (AD3) board. Signal modulation and recording are performed on host computers tethered to the ASV and ROV via AD2 and AD3 devices. We capture signal strength, represented by the induced voltage on the ME receiving array, as a function of distance. This data was recorded using Digilent’s WaveForms software on the host computers.
4 System Integration & Experimental Setup
Our experimental analyses include (i) evaluation of the BlueME system’s magnetoelectric (ME) antenna communication characteristics, and (ii) performance assessment through robotics field trials. First, we validate the proposed antenna’s functionality using empirical data and simulation results, providing a basis for further investigation into optimal modulation techniques. We then integrate two BlueME systems on an autonomous surface vehicle (ASV) and an underwater remotely operated vehicle (ROV) to evaluate communication range and robustness, as illustrated in Fig. 6. To this end, we conduct a series of field experimental trials in open water to demonstrate the feasibility of the proposed BlueME system for robot-to-robot communication.
4.1 Setup: Measuring Antenna Impedance
Before assembling the full antenna arrays, each ME antenna was tested individually to determine its resonant frequency and impedance characteristics. The objective was to verify that the empirical measurements matched our simulation results and to assess the combined response of the antennas when configured in arrays. Each antenna was placed in a shallow plastic vat filled with oil to replicate the dielectric environment in the enclosure. The antennas were then positioned within the array frame using bias magnets to ensure that the magnetic field conditions matched the assembled array, as their resonant response depends on the magnetic bias levels, which can vary due to the arrangement of magnets in the array frame.
(a) Freshwater experiments (lake trials)
(b) Saltwater experiments (ocean trials)
An impedance sweep from to was performed using a DAD3 board connected to an impedance measurement attachment and a BNC adapter. In these tests, the antennas were connected via short wires to a terminal and BNC oscilloscope probes. The impedance magnitude and phase were recorded for each antenna individually.
After individual tests, the antennas were wired together to form the transmitter and receiver arrays. The transmitter antennas were connected in parallel, while the receiver antennas were connected in series. The arrays were then tested as a whole, with impedance measurements taken under the same conditions as the individual antennas. The goal of this setup was to fabricate the ME antennas with sufficiently tight tolerances to validate that wiring the antennas in parallel or series effectively increases bandwidth and/or resonant response [62].
4.2 Setup: Underwater Communication Range
Lake Trials: We first conducted proof-of-concept experiments by evaluating the impedance of the proposed system in a water tank measuring m (L) m (W) m (D). The range tests were conducted using two ME antenna arrays in a freshwater lake. The receiver array was mounted on an underwater ROV (BlueROV2) and positioned at a depth of approximately below the surface, whereas the transmitter array was mounted on an ASV (BlueBoat) with the transmitter submerged approximately below the surface. The transmitter array was driven by a Class-D amplifier powered by a supply, which was generated using a boost converter connected to a LiFePO4 battery. The input signal to the amplifier varied from to . A small monitoring board was placed in line with the amplifier output, consisting of current sense resistors and an AC-coupled : voltage divider to measure the current through and voltage across the antenna array.
The receiver array (on the ROV) was connected to an AlphaLab LNA10 low-noise amplifier (LNA) set to gain with a low-pass filter cutoff at approximately to mitigate aliasing effects. The output of the LNA was connected to a DAD3 board for data acquisition at a sampling rate of . The LNA and DAD3 were housed in a waterproof enclosure mounted on the ROV. Both transmitter and receiver arrays were connected using RG-6 coaxial cables with BNC compression fittings.
For the range tests, the transmitter array conducted frequency sweeps from to at varying drive strengths. The receiver array collected data continuously, using the FFT peak hold feature of the WaveForms software to record the received signal amplitude across the frequency range. Initial tests were performed in freshwater at distances of , , , and . The ASV’s position during each trial was estimated via GPS with an accuracy of approximately and used to determine the transmitter–receiver separation (see Fig. 6). The maximum test distance was set to based on the available test area at the lake.
Ocean Trials: For more comprehensive and continuous testing, the experiments were repeated in saltwater off the west coast of Florida (Fig. 9). The Gulf test sites ranged from depths of approximately – (– m); range measurements were performed for horizontal distances. As shown in Fig. 6, the Tx (mounted on ASV) drifted continuously away, with the Tx assembly suspended 3 m below the water surface, while the Rx (mounted on ROV) operated near the ship at depths of -. GPS data from the ASV and the ship were used to determine the TX–RX separation, which was measured up to .
5 Performance Evaluation: Freshwater
5.1 Impedance Measurement Results
The impedance measurements of the individual antennas are shown in Fig. 7; the results match our simulations and proof-of-concept experiments, thus validating the proposed design. We observe that the parallel configuration of transmitter antennas resulted in an impedance close to the parallel sum of the individual antenna impedances within the target frequency range. Similarly, the series configuration of the receiver antennas yielded an impedance close to the sum of the individual impedances. These results demonstrate that wiring the antennas in parallel or series effectively adjusts the overall impedance and resonant response, which can improve bandwidth and signal strength; see Fig. 8.
Furthermore, some antennas exhibited smaller resonant peaks near the primary resonant frequency of –. We hypothesize that this is due to minor variations in the fabrication process or internal micro-cracks. Notably, certain antennas displayed approximately one order of magnitude lower impedance at resonance and higher quality factors than average, indicating stronger resonant responses. Conversely, some antennas had approximately one order of magnitude higher resonant impedance and lower quality factors than average, suggesting weaker resonant responses but wider bandwidths. These variations highlight the importance of consistent fabrication processes and the potential for optimization through mass manufacturing and binning.
5.2 Freshwater Communication Performance
For the range tests, we initially anticipated that reactive near-field behavior would dominate the antenna response, resulting in rapid signal attenuation with distance. However, the measured results exhibited a slower decay rate than expected from purely reactive near-field coupling, with signals observed from to . These observations indicate that part of the transmitted energy propagated beyond the reactive near field, possibly through radiative or interface-guided modes along the water–air or water–seabed boundaries [63]. The maximum test distance in freshwater was set to based on the available test area; signal detection beyond this range was subsequently confirmed in ocean trials (see Sec. 6).
This performance can be attributed to several factors. First, the electromagnetic wavelength in water at is approximately , much shorter than in air due to water’s higher permittivity and conductivity. The wavelength in a conductive medium is given by , where is the phase constant. In conductive media, is influenced by permittivity, permeability, conductivity, and angular frequency. The reduced wavelength improves the radiation efficiency of electrically small antennas. Additionally, the ME antennas, when configured in arrays, are within each other’s near fields, effectively forming a larger equivalent magnetic dipole with increased radiation [11]. Antenna misalignment had a more pronounced effect at closer ranges. This can be attributed to near-field interactions causing destructive interference at close range and/or signal propagation along more diverse paths at greater distances.
Moreover, we found that the rate of attenuation decreased with increasing distance. This may be due to electromagnetic waves propagating along paths with lower attenuation (e.g., the lake bed, surface wave) [64] or due to reflections at shallow surfaces [15]. Unlike acoustic propagation, electromagnetic waves’ significantly higher propagation velocity in water results in wavelengths of orders of magnitude longer than acoustic waves at comparable frequencies. Consequently, typical path length differences produce smaller phase differences between multipath components, rendering interference effects negligible. Additionally, reflections from obstacles and natural features are minimal since objects of size comparable to or smaller than the electromagnetic wavelength tend to produce weak scattering, with incident waves primarily diffracting around or penetrating through such objects rather than producing strong reflections.
5.3 Analyzing the Theory of Operation
The performance characteristics of ME antenna arrays submerged underwater are not extensively explored in the literature. As such, we make several observations in our analysis.
Wavelength Changes. The antenna’s observed rate of signal drop-off appears to change across distances. This may be explained by the change in electromagnetic wavelength from air to fully submerged in water, which is also a more conductive medium. To this end, we calculate the expected wavelength of an electromagnetic plane wave in an underwater environment. As shown by [14], the electromagnetic propagation constant can be calculated as:
where is the water medium’s permeability, denotes conductivity, and is the angular frequency. Given that the displacement current may be neglected at sufficiently low frequencies [14], can be expressed as:
where and are the attenuation and phase constants, respectively, and in this case. Subsequently, the expression for wavelength in terms of and is:
The specific conductivity of Lake Wauburg in Gainesville, Florida (N, W) is approximately as measured over a 16-year averaging period [65], which equals . Using the formula above, we calculate the wavelength of an electromagnetic wave to be between approximately and for frequencies from to , respectively. Thus, the underlying frequency has a wavelength of approximately , compared to approximately in air.
Far-Field Effects. For electrically small antennas, the reactive near-field region extends to from the antenna, which is roughly in our case. The radiative near-field extends to approximately one wavelength away [62]. The transition zone between the near-field and far-field regions lies between and . Therefore, the far-field region of the antenna begins from and beyond.
Radiation Resistance. The radiation resistance characterizes the efficiency of an antenna in radiating power [62, 66]; it is calculated by the following expression:
where is the wave impedance of the medium, is the phase constant, is the antenna length, and is the wavelength. In water, the wave impedance is , where is the free-space impedance and is the relative permittivity of water. Thus, becomes:
Assuming that and all other constants remain identical between air and water, we find
Although the relative permittivity of water is significantly higher than air (:), the substantial decrease in wavelength (from in air to in water) results in a net increase in radiation resistance per antenna by a factor of approximately . This leads to an increase in the total radiated power per antenna by a similar factor.
Single Antenna vs Antenna Array. Increasing number of antennas on the transmitter side ( to ) produces a quadratic increase in net radiated power, as shown by Dong et al. [11]:
Since the constants remain the same across array configurations and is linearly proportional to the number of ME antenna elements in the transmitter array , we have
Thus, an increase in total radiated power for a array on the transmitter side is a factor of , assuming negligible mutual coupling losses.
Increasing the number of antennas on the receiver side also increases the receiver sensitivity linearly, assuming all antennas are identical, i.e., a factor of improvement in the overall link budget. Consequently, using an array of antennas underwater compared to a single antenna in air theoretically yields an overall link budget improvement by a factor of , which corresponds to an increase of approximately dB – a factor not specifically discussed in the literature [11, 48, 29, 34, 35]. In practice, however, resonant frequency mismatches and increased mutual coupling at closer antenna spacings and higher drive strengths will degrade phase coherence within the transmitter array [62], reducing the actual achievable improvement from the array configuration.
5.4 Power Footprint and Robustness
The proposed BlueME system consumes minimal power, with optimal performance achieved at approximately of power consumption (see Fig. 10). Nonlinear effects such as spring-softening in the piezoelectric materials and spring-hardening in the Metglas layers start to become significant at higher drive strengths, leading to shifts and distortions in the resonant frequency with decreased quality factors [67]. In Fig. 10, increased drive power shifts the resonance peak to lower frequencies and distorts the response shape. Shown earlier in Fig. 8, the received signals at various transmitter power levels and distances, a downward frequency shift is also observed. However, since the receiver experiences only the coupled power from the transmitter, these power levels are much lower, resulting in minimal nonlinear distortion at the receiver. Thus, the observed frequency shift primarily reflects the transmitter’s nonlinear behavior combined with the receiver’s nearly constant resonant response. These effects can be exploited to adjust for resonance mismatches between RX/TX arrays by varying the drive strength. We also observed that the antenna performance was not significantly affected by obstacles, water turbidity, or multipath propagation, which are major challenges for traditional acoustic and optical modalities.
5.5 Channel Capacity
The development of an optimal modulation scheme is beyond the scope of this paper; however, we can derive theoretical channel limits from our measurements using the Shannon-Hartley theorem [68]. For a simple channel with additive white Gaussian noise, the Shannon-Hartley theorem gives the channel capacity as:
where is the channel bandwidth, is the received signal power, and is the noise power. The signal and noise powers can be calculated from their respective RMS voltages squared. Based on measurements of the LNA noise floor, the noise spectra within our frequency range of interest (31 kHz to 41 kHz) can be approximated as Gaussian. For our frequency-selective underwater channel, we must consider the varying SNR across different frequencies. This leads to a more general form of the capacity formula:
Here, and represent the frequency-dependent signal and noise power spectral densities. Since our measurements are taken at discrete frequency points (shown in Fig. 8), we can express this as:
where is the frequency index, is the total number of frequency points, and is the frequency resolution of our measurements (). Using this discretized form, we calculate the cumulative theoretical channel capacity from our signal strength measurements, with results shown in Fig. 11.
6 Performance Evaluation: Saltwater
Antenna performance was further evaluated in saltwater trials conducted off the west coast of Florida along the Gulf (see Fig. 9). The primary objective was to examine the effect of higher conductivity () on achievable communication range. The experimental setup mirrored that of the freshwater tests, with all equipment housed in watertight enclosures and mounted on the BlueBoat (Tx) and BlueROV2 (Rx). Minor modifications included the integration of a LattePanda single-board computer for autonomous control and the replacement of the original Class-D amplifier with a smaller unit to accommodate spatial constraints. During testing, all systems were powered by onboard batteries and operated fully submerged, except for an Ethernet tether used to stream real-time data from the receiver. For the saltwater trials, the LNA gain was reduced to 40 dB from 60 dB in the freshwater tests to prevent saturation, which lowered the measured noise floor but did not measurably affect the overall SNR.
Autonomous frequency sweeps between to were conducted at one-minute intervals by the transmitter. The BlueBoat carrying the transmitter was allowed to drift freely, with its position tracked using GPS. The receiver was positioned approximately from a stationary pontoon boat and recorded peak-hold FFT measurements each time new GPS coordinates were logged.
An initial transmitter power sweep (shown in Fig. 16) confirmed negligible differences in antenna characteristics relative to the freshwater experiments (shown earlier in Fig. 10), validating the consistency of power characteristics across test conditions. Subsequent experiments involved transmitter sweeps at fixed distances under varying input power levels, as well as tests at fixed transmitter power with varying transmitter–receiver separation. From these measurements, the received SNR was computed, and the corresponding spectral characteristics were analyzed, as presented in Figs. 13 and 14.
6.1 Wavelength and Field Regions
The near-field, transition, and far-field boundaries were estimated using the same method applied in the freshwater analysis. Conductivity at the site was and the water temperature ranged from 20–25∘C [69]. For a center frequency of and seawater conductivity of , the wavelength is calculated as . The corresponding theoretical boundaries for the antenna regions are defined as:
Reactive near-field ():
Radiative near-field ():
Transition region ():
Far-field ():
The measured SNR at distances below approximately exhibited a steeper decay than the nominal – dependence typically associated with reactive or radiative near-field behavior. At larger separations (), the attenuation trend approached an approximate dependence, suggesting that signal propagation may have occurred along paths not confined to the bulk medium, such as along the water–air or water–seabed interfaces [63, 70]. A localized deviation from this overall trend was observed at a transmitter–receiver spacing of (Fig. 13), where the peak received signal level was approximately higher than that measured at . A similar deviation was noted at , where the peak signal level exceeded that at by roughly . A reduction in the measured noise floor was observed during the test, caused by a drop in the LNA supply voltage as its battery approached full discharge. All measurements were performed under identical experimental conditions, with the submerged transmitter naturally drifting and rotating during the tests.
6.2 Radiation Resistance
The same relationship used in the freshwater case can be extended to saltwater by considering the higher conductivity of the medium. Starting from the general dependence
and substituting for , the radiation resistance becomes directly dependent on the conductivity of the surrounding medium. Assuming the frequency, permeability, and relative permittivity remain constant between freshwater and saltwater, this simplifies to
Substituting representative conductivities for freshwater () and saltwater () gives the expected ratio as
This result follows directly from the same scaling principles but reflects the influence of saltwater’s greater conductivity on the radiation resistance of the antenna.
6.3 Power Consumption and Nonlinear Effects
The power consumption characteristics measured during the saltwater tests (Fig. 16) were consistent with those obtained earlier in freshwater (Fig. 10). The same antenna assemblies were used in both experiments, and the response curves exhibited comparable shapes and resonance features. Nonlinear behavior became evident at higher drive levels, with the resonance peak broadening and shifting asymmetrically toward lower frequencies. In the received-signal measurements (Fig. 14), the overall signal amplitude increased with drive strength up to approximately , beyond which additional input power did not produce proportionally higher received levels. A deviation from this trend was observed at an input of , where the peak received signal level exceeded that measured at and .
6.4 Modulation and Practical Considerations
When an ME antenna is driven at high amplitudes, its effective stiffness becomes amplitude dependent, causing the resonance to deviate from the Lorentzian form and exhibit Duffing-type nonlinear behavior [71, 72, 73, 74]. When a modulated signal is applied, these nonlinearities generate harmonic and intermodulation distortion [75, 76]. Amplitude-varying modulation formats are particularly susceptible to spectral regrowth and amplitude-to-phase conversion. Constant-envelope, noncoherent formats such as MSK, CPFSK, and BFSK are therefore preferred for communication; alternative approaches such as direct antenna modulation (DAM) and nonlinear antenna modulation (NAM) have also been explored to intentionally utilize the nonlinear response within the modulation process [77, 78].
ME antennas inherently support multiple resonant modes beyond the fundamental, corresponding to higher-order mechanical and electromechanical harmonics of the piezoelectric element [79]. In the present system, additional resonances appear above approximately 70 kHz and extend to several hundred kHz. Operation outside the fundamental band can excite these higher modes if the transmit spectrum includes harmonic components or spectral leakage from the modulation process. Maintaining spectral confinement helps suppress spurious excitation of these bands, although such measures are unnecessary when operation is restricted to a single frequency range.
The radiated field of an ME antenna originates from the time-varying magnetic flux density produced by magnetostrictive motion, which undergoes a phase transition of nearly through resonance [80]. This frequency-dependent phase transition produces a corresponding frequency-dependent group delay in the radiated signal. The overall group delay of the communication link reflects the combined phase responses of the Tx and Rx antennas. For narrowband operation (), this delay remains effectively constant and waveform distortion is negligible, whereas for wider bandwidths (), phase nonuniformity can distort the transmitted signal unless characterized and compensated at both the transmitter and receiver. Given these antenna characteristics, the desirable properties of a suitable modulation scheme are: (i) constant envelope, to minimize nonlinear conversion; (ii) spectral compactness, to avoid excitation of adjacent resonant modes; and (iii) resilience to phase dispersion, or compatibility with digital equalization schemes.
Binary frequency shift keying (BFSK) is a representative modulation scheme satisfying these requirements when operation is confined to the fundamental resonance band. In BFSK, binary symbols are conveyed by switching the carrier between two frequencies, and , ideally within the resonance bandwidth. Because the envelope remains constant and the frequency deviation can be selected to stay within the array’s linear operating range, BFSK avoids AM-PM conversion and minimizes distortion. Under small-signal conditions where each antenna element operates well below its nonlinearity threshold, theoretical performance can be analyzed using the standard BER–SNR relationship.
Noncoherent BFSK compares the received energy in two frequency bands, so the decision statistic depends only on signal magnitude and is insensitive to carrier phase. This phase independence makes it well-suited to ME antennas, where the radiated and detected phase varies across the resonance bandwidth unless explicitly characterized and compensated at the transmitter and receiver. Tone separation can therefore span a large fraction of the resonant band without affecting detection. The corresponding BER in additive white Gaussian noise (AWGN) [81, 82, 83] is computed as:
The maximum theoretical bit rate for BFSK is determined by the minimum frequency separation required to maintain symbol orthogonality. For noncoherent detection, orthogonality is achieved when the two tones differ by an integer number of cycles over one bit interval , such that . More generally, the energy-per-bit to noise-density ratio can be expressed in terms of the SNR, the bit rate , and the tone spacing as
where is taken as the minimum SNR of the two selected tones. Substituting this into the noncoherent BFSK bit-error rate equation can be reduced as follows:
Given the measured SNR as a function of frequency for various transmission distances (Fig. 13 and Fig. 8), the maximum achievable bit rate for a desired can be estimated by selecting two tones near the resonance frequency that satisfy the BFSK tone-spacing condition and exhibit approximately equal SNR. For a chosen BER threshold, the maximum tone separation that satisfies this criterion can be used to determine , which can then be mapped to a range via the corresponding SNR–distance relationship. This analysis, assuming an uncoded channel, is illustrated in Fig. 15.
7 Analysis & Discussion
7.1 Comparison with Other Underwater Communication Modalities
The proposed BlueME system addresses several underwater communication challenges that commonly affect acoustic, optical, and higher-frequency radio approaches. We now provide a theoretical grounding for the observed properties, cost implications, and operational trade-offs; the summary is presented in Table 1. The tabulated data rates, costs, and ranges are derived from typical systems available in academia and industry. The listed values represent approximate order-of-magnitude maxima commonly reported for each method, serving as representative baselines for typical devices and operating conditions.
| System | Advantages | Disadvantages |
Examples | |||||||||||||
| Acoustic |
10 b/s to 10 kb/s | 100 m to 100 km | 1k to 10k | [84, 85, 86, 87] | |||||||||||||
| Optical |
1 Mb/s to 1 Gb/s | 10 m to 100 m | 1k to 10k | [88, 89, 90, 91, 92] | |||||||||||||
100 kb/s to 1 Mb/s | 1 m to 100 m | 100 to 1k | [93, 94, 95, 96, 97] | |||||||||||||
100 kb/s to 1 Mb/s | 100 cm to 1 m | 10 to 100 | [98, 99, 100, 101, 102, 103] | |||||||||||||
100 kb/s to 1 Mb/s | 1 m to 10 m | 100 to 1k | [104, 105, 106, 107] | |||||||||||||
10 m to 100 m | 100 to 1k | [108, 109] | |||||||||||||
• Affected by water conductivity | 1 kb/s to 100 kb/s | 10 m to 1 km | 10 to 100 |
Acoustic links have been the preferred standard for underwater applications, offering long-range propagation up to km with relatively modest power consumption [84, 85]. Sound waves couple efficiently into the water and travel long distances with lower attenuation than electromagnetic waves [86, 87]. However, acoustic signals experience substantial multipath, Doppler shifts, and boundary reflections, especially in dynamic marine environments. These effects limit typical data throughput to the order of b/s– kb/s and also raise concerns about potential harm to nearby humans and marine fauna if high-intensity acoustic sources are used.
Optical methods such as laser or LED-based systems can provide data rates in excess of Gb/s with low latency [88, 90]. However, even moderate turbidity significantly degrades optical transmissions, restricting practical operational ranges to a few tens of meters. Narrow-beam alignment requirements further complicate deployment in many natural waters where visibility can be variable [89, 91, 92].
Magnetic-field communication relies on near-field inductive coupling via magnetic coils. These methods are often employed for through-the-earth or short-to-medium-range underwater links and can operate over 1 m to 100 m with data rates up to about 1 Mb/s [93, 94]. Magnetic fields typically penetrate obstacles effectively and offer high reliability where direct line-of-sight may be obstructed. However, large inductors or ferrite materials are often required to generate sufficient field strength, leading to heavier hardware requirements. Accurate alignment between transmitter and receiver coils is also necessary to maximize coupling efficiency [95, 96, 97]. Typical costs for such systems are on the order of $100 to $1 k USD.
Electric-field communication relies on generating and sensing electric fields in conductive media using capacitive or galvanic coupling [98, 99, 101, 102, 103]. It can be implemented with lightweight and inexpensive electrodes, potentially supporting simultaneous power transfer and data communication. Still, the typical range is limited to about m and there are important safety considerations when using higher voltages around marine life or humans [100].
Standard RF methods that operate at medium to high-frequency (MF/HF/VHF) exhibit higher data rates, typically on the order of 1 Mb/s. However, the attenuation in conductive media is much higher, which limits transmission distances to about - m [104, 105, 106, 107]. These higher-frequency systems may also require specialized antenna calibration to address de-tuning effects caused by changing dielectric properties underwater. Recent demonstrations of surface-electromagnetic-wave (SEW)-based RF communication have reported ranges exceeding 100 m and bandwidths of several megahertz by utilizing near-surface propagation along the water-air interface [108, 109]. In these implementations, electromagnetic energy is guided along the boundary layer rather than through the bulk medium, resulting in lower effective attenuation compared to volumetric propagation. This mode of operation represents a distinct configuration from conventional submerged RF systems and is therefore listed as a separate class in Table 1.
In comparison, our ME-based RF systems in the VLF/LF, or medium-frequency (MF) range rely on magnetostrictive-piezoelectric hetero-structures that convert magnetic excitation into electrical signals, and vice versa [11]. One of their primary advantages is immunity to turbidity and obstructions, which often hinder optical and acoustic communication. The longer electromagnetic wavelengths (on the order of tens to hundreds of meters in freshwater) reduce phase interference and Doppler effects, thereby mitigating multipath fading. Although water conductivity does lead to attenuation at increasing distances, the experimentally validated propagation ranges remain practically useful. Fabrication costs for ME-based antennas at the scales used in our experiments can be on the order of a few USD per element, making them relatively cost-effective. Power requirements are low, and the environmental impact is minimal, as these systems generate negligible acoustic noise and avoid light emissions that could disturb marine ecosystems.
Overall, the comparison in Table 1 indicates that no single modality fully addresses all underwater communication requirements. Acoustic systems excel at long-range transmission but are constrained by low data rates. While higher-frequency RF and optical methods can provide larger bandwidths, they operate at severely limited ranges in conductive or turbid waterbody conditions. In contrast, ME-based RF offers a promising balance of moderate-range coverage, moderate data rates, low power consumption, and low risk to marine life. This trade-off analysis suggests that the BlueME system can occupy an attractive niche in scenarios where turbidity or multipath might compromise other modalities while remaining cost-effective and environmentally benign.
7.2 Future Potential and Practicalities
In future implementations, it may be desirable to implement custom FSK schemes for digital communication. FSK will potentially minimize the impact of amplitude-based nonlinear distortions, allowing for reliable data transmission even in the presence of nonlinearities at higher drive strengths. The ability to pack multiple antennas into arrays provides opportunities for scaling communication capabilities. Bandwidth and power can be increased by adding more antennas, and arrays can be optimized for specific operational frequencies or bandwidth requirements.
Overall, the proposed BlueME system is an attractive option for real-time underwater sensor networks and robotics applications. We envision that it could be used for close-range robot-to-robot communication as well as for coordinated planning, relative localization, and control by operating at higher frequencies with shorter wavelengths.
8 CONCLUSION
In this paper, we have presented the design, implementation, and experimental validation of a novel ME antenna array system, BlueME, for underwater communication in robotics applications. The BlueME system offers an alternative to traditional acoustic and optical communication methods, addressing challenges in range, efficiency, robustness, and scalability. In testing, the BlueME system achieved reliable communication at distances of over with a power consumption of only in fresh water and over in sea water with a consumption of , and indicating its suitability for low-power platforms, such as mobile robots and standalone sensors. Our results also demonstrate that the system remains unaffected by underwater noise, turbidity, and multipath interference.
Future work will focus on optimizing the antenna fabrication process to ensure consistent performance and on tailored modulation schemes that extend the system’s capabilities. This includes enabling applications in multi-robot cooperative tasks such as localization, navigation, and mapping, where reliable, low-power underwater communication is essential.
ACKNOWLEDGMENT
This research is supported in part by the U.S. National Science Foundation (NSF) grants # and #; and the University of Florida Research grant #.